Laborjournal vom 2003-04-22 bis …

Erika Ehrgeizig 11. August 2003

Index

Expression von Hagridin in E. coli, 3

Transformation mit HagIV-Plasmid, 3

Herstellung von A, 5 Herstellung von B, 7 Herstellung von gereinigtem Irgendwas,

5

Lyophilisieren von A, 7 Lyophilisieren von B, 7

Präparation von RNase-A-S-Peptid, 1–2,

5 Subtilisin-Verdau, 1–2, 5 Index iv

Inhaltsverzeichnis

Dienstag, 22. April 2003 1

1 Präparation von RNase-A-S-Peptid . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Subtilisin-Verdau . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Expression von Hagridin in E. coli . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Transformation mit HagIV-Plasmid . . . . . . . . . . . . . . . . . . 3

Mittwoch, 23. April 2003 5

1 Präparation von RNase-A-S-Peptid . . . . . . . . . . . . . . . . . . . . . . 5

1.1 Subtilisin-Verdau . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2 Herstellung von gereinigtem Irgendwas . . . . . . . . . . . . . . . . . . . . 5 3 Herstellung von A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Donnerstag, 24. April 2003 7 1 Herstellung von B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2 Lyophilisieren von A und B . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Inhaltsverzeichnis vi

Dienstag, 22. April 2003

1 Präparation von RNase-A-S-Peptid

1.1 Behandlung mit Subtilisin

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

Dienstag, 22. April 2003

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges. A rather lengthy description which is just here to use some space, so that one can see the effect of the index ranges.

Abbildung 1: Ein sehr interessantes SDS-Gel

2 Expression von Hagridin in E. coli

2.1 Transformation von hogwarts-auxotrophen E. coli mit dem Plasmid HagIV

Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Some shorter Text, the cells will only grow if you’re not a muggle… Dienstag, 22. April 2003 4

Mittwoch, 23. April 2003

1 Präparation von RNase-A-S-Peptid 1.1 Verdau mit der neuen Charge Subtilisin 2 Herstellung von gereinigtem Irgendwas 3 Herstellung von A

A wurde nach der Vorschrift in Müller et al. hergestellt und vorläufig im Eisschrank gelagert.

Mittwoch, 23. April 2003 6

Donnerstag, 24. April 2003

1 Herstellung von B

B wurde nach der Vorschrift von Meier und Schulze hergestellt.

2 Lyophilisieren von A und B